Kate's father is 32 years older than Kate is. In 5 years, he will be five times older. How old is Kate?
Answer by stanbon(75887)

Kate's father is 32 years older than Kate is. In 5 years, he will be five times older. How old is Kate?
---------------
Equations::
f = k + 32
f+5 = 5(k+5)
------------------------
Substitute for "f" and solve for "k":
k+32 + 5 = 5k+25
k + 37 = 5k + 25
4k = 12
Kate's age now = 3 yrs
-----------
Cheers,
Stan H.